OpenStudy (anonymous):

find the parametric equation: given a curve with the vertex (1,-9) and passes through (0,0) and (2,0). The answer should have x and y as functions of t. Help? i don't know how to start?

OpenStudy (phi):

it sounds like they want a parabola y= 9x(x-2) you could let x=t and y= 9t^2 -18t

OpenStudy (anonymous):

@phi i tried to find the slop between those two points given and got zero.. so i plugged that m=0 in y-y0=m(x-xo) i plugged the vertex coordinates into yo and Xo and got y=-9x.. then i stated that y=t and solved for x and got x=t/-9... does that sound reasonable?

OpenStudy (phi):

you are trying to fit a line through the 3 points. if you plot them, they look something like this: